# Automatic Pruning via Structured Lasso with Class-wise Information

### v1

@@ 5 Conclusion
We propose two novel automatic pruning methods that both achieve state-of-the-art, sGLP-IB and sTLP-IB, which leverage information bottleneck theory and structured lasso to perform pruning using penalties that construct subgroup and subtree structures based on class-wise statistical information. Additionally, we present an adaptive approach that utilizes Gram matrix to accelerate the process and employs proximal gradient descent to further ensure convergence and polynomial time complexity. Through extensive experiments and ablation studies, we demonstrate the potential of our methods, both theoretically and experimentally, to effectively utilize class-wise information to aid pruning tasks, compared with numerous baselines. Exploring the better method to capture the class-wise information is our future direction.
Appendix A Experiments
A.1 Dataset Information
@@
We used three widely adopted datasets to evaluate our proposed methods against baselines: CIFAR10 Krizhevsky [2009], CIFAR100 Krizhevsky [2009] and Tiny ImageNet Deng et al. [2009].
@@
A.2 Time Complexity Discussion
For sGLP-IB: the time complexity of graph-structured lasso could be found in the paper Chen et al. [2010].
For sTLP-IB: since a tree associated with $out_{i}$ responses can have at most $12out_{i}-$ 1 nodes, which constrains the $|G_{v}|$. It is computationally efficient and spatially economical to run sTLP-IB. We employ a smoothing proximal gradient method that is originally developed for structured-sparsity-inducing penalties. By using the efficient method, the convergence rate of the algorithm is $O(\frac{1}{\epsilon})$, given the desired accuracy $\epsilon$ and the time complexity per iteration of the smoothing proximal gradient is $O(in_{i}^{2}out_{i}+in_{i}\sum{v\in V}|G_{v}|)$. Thus the overall complexity for our method is $O(\frac{1}{\epsilon}\times(in_{i}^{2}out_{i}+in_{i}\sum{v\in V}|G_{v}|))$.
A.3 Details of Pruning on ResNet and GoogleNet
For ResNet: due to ResNet’s architecture featuring shortcuts, when pruning ResNet variants, we initially ignore the shortcuts. We perform pruning layer by layer, and then prune this shortcut layer corresponding to the pruned channels of its related layers inside the residual block when we approach the end of this residual block. This pruning approach ensures the integrity and functionality of the shortcut connections while effectively reducing the network’s complexity.
@@ A.3 Details of Pruning on ResNet and GoogleNet
Table 8: Pruning results of VGG-16 on CIFAR-100 dataset.
A.4 Experiments with CIFAR-100 Dataset
We compare sGLP-IB and sTLP-IB with CPMC Yan et al. [2021], PGMPF Cai et al. [2022], CPGMI Lee et al. [2020], PGMPF-SFP Cai et al. [2022], and APIB Guo et al. [2023] on the CIFAR-100 dataset using VGGNet-16. Our methods, sGLP-IB and sTLP-IB, achieve the highest accuracy with the largest pruning ratio (55% for FLOPs and 71% for parameters), reaching 73.96% and 73.99%, respectively, which are 0.16% and 0.19% higher than the base model. This comparison with these state-of-the-art methods highlights the effectiveness of our approach, considering Top-1 accuracy, FLOPs pruned ratio, and parameter reduction ratio metrics. Combined with results from experiments on CIFAR-10 and ImageNet, it demonstrates that our methods can be extended to various real-world datasets.
@@ A.4 Experiments with CIFAR-100 Dataset
A.5 Experiments with Pruning Ratios for FLOPs
Before the pruning ratio exceeds 50%, we observe that the parameter pruned ratio is also below 60%, and the performance of sGLP-IB and sTLP-IB is nearly identical. It is noteworthy, however, that sTLP-IB prunes more parameters, indicating that it retains fewer parameters while maintaining the same FLOPs. This suggests that sTLP-IB is more effective at focusing on important filters due to its superior ability to model class-wise information. When the pruning ratio exceeds 70%, the accuracy of both methods begins to drop. At a ratio of 80%, the accuracy of our methods drops by 2-3% due to excessive parameter reduction (94%), though sTLP-IB still remains more robust than sGLP-IB.

### v2

@@ Conclusion
We propose two novel automatic pruning methods that both achieve state-of-the-art, sGLP-IB and sTLP-IB, which leverage information bottleneck theory and structured lasso to perform pruning using penalties that construct subgroup and subtree structures based on class-wise statistical information. Additionally, we present an adaptive approach that utilizes Gram matrix to accelerate the process and employs proximal gradient descent to further ensure convergence and polynomial time complexity. Through extensive experiments and ablation studies, we demonstrate the potential of our methods—both theoretically and experimentally—to effectively leverage structured class-wise information for pruning tasks and facilitate deployment on resource-constrained devices, compared with numerous baselines.
Appendix A Artifact Evulation
In the README file of our submitted code repo, you will find all the necessary bash scripts to reproduce main tables and figures from our experiments. The hyperparameter details are all in our code.
Appendix B Time Complexity and Convergence Discussions
Time Complexity Discussions
We employ a smoothing proximal gradient method that is originally developed for structured-sparsity-inducing penalties. By using the efficient method, the convergence rate of the algorithm is $O(\frac{1}{\epsilon})$
For sGLP-IB: the time complexity per iteration of the smoothing proximal gradient for sGLP-IB is $O(in_{i}^{2}out_{i}+in_{i}|E|)$. Thus, the overall time complexity for sGLP-IB is $O(\frac{1}{\epsilon}\times(in_{i}^{2}out_{i}+in_{i}|E|))$.
For sTLP-IB: since a tree associated with $out_{i}$ responses can have at most $2out_{i}-1$ nodes, which constrains the $|G_{v}|$. It is computationally efficient and spatially economical to run sTLP-IB. The time complexity per iteration of the smoothing proximal gradient is $O(in_{i}^{2}out_{i}+in_{i}\sum{v\in V}|G_{v}|)$. Given the desired accuracy $\epsilon$, the overall complexity for our method is $O(\frac{1}{\epsilon}\times(in_{i}^{2}out_{i}+in_{i}\sum{v\in V}|G_{v}|))$.
Convergence Discussions
We rewrite $bs\times bs$ as $N$, $\bm{X}_{i}^{g}$ as $\bm{X}$, $in_{i}$ as $J$, $out_{i}$ as $K$, $\bm{X}_{i+1}^{g}$ as $\bm{Y}$. Then we will get the original form of Lasso as the following:
[…]
The edge set $E$ for sGLP-IB is generated similarly to graph- structured lasso; the $G_{v}$ in sTLP-IB is also generated following tree-guided lasso. Thus, $\Phi(\bm{\beta})$ follows the original forms of the previous two lasso variants. Based on these, our methods also follow the proximal gradient descent framework (Chen et al. 2012; Yuan 2024) and are applied to optimization problems of the same format. The format is as follows:
$\min\limits_{x}f(x)+g(x)$ (9)
[…]
Thus, our methods, sGLP-IB and sTLP-IB, possess the property of convergence
Figure 3: Our example experimental prototype for resource-constrained devices.
Appendix C Experiments
Adaption for Pruning ResNet and GoogleNet
As architectures of ResNet and GoogleNet differ from conventional CNNs, we make adaptations to the pruning process accordingly.
For ResNet: due to ResNet’s architecture featuring shortcuts, when pruning ResNet variants, we initially ignore the shortcuts. We perform pruning layer by layer, and then prune this shortcut layer corresponding to the pruned channels of its related layers inside the residual block when we approach the end of this residual block. This pruning approach ensures the integrity and functionality of the shortcut connections while effectively reducing the overall network’s complexity.
Dataset Information
@@ Appendix C Experiments
Datasets: We used three widely adopted datasets to evaluate our proposed methods against baselines: CIFAR10 (Krizhevsky 2009), CIFAR100 (Krizhevsky 2009) and ImageNet (Deng et al. 2009).
@@
Table 8: Pruning results of VGG-16 on CIFAR-100 dataset.
@@
Experiments with CIFAR-100 Dataset
We compare sGLP-IB and sTLP-IB with current methods in Table 8 on the CIFAR-100 dataset using VGGNet-16. Our methods, sGLP-IB and sTLP-IB, achieve the highest accuracy with the largest pruning ratio (55% for FLOPs and 71% for parameters), reaching 73.96% and 73.99%, respectively, which are 0.16% and 0.19% higher than the base model. This comparison with these state-of-the-art methods highlights the effectiveness of our approach, considering Top-1 accuracy, FLOPs pruned ratio, and parameter reduction ratio metrics. Combined with results from experiments on CIFAR-10 and ImageNet, it demonstrates that our methods can be extended to various real-world datasets.
@@
Experiments with Pruning Ratios for FLOPs
In Table 9, before the pruning ratio exceeds 50%, we observe that the parameter pruned ratio is also below 60%, and the performance of sGLP-IB and sTLP-IB is nearly identical. It is noteworthy, however, that sTLP-IB prunes more parameters, indicating that it retains fewer parameters while maintaining the same FLOPs. This suggests that sTLP-IB is more effective at focusing on important filters due to its superior ability to model structured class-wise information. When the pruning ratio exceeds 70%, the accuracy of both methods begins to drop. At a ratio of 80%, the accuracy of our methods drops by 2-3% due to excessive parameter reduction (94%), though sTLP-IB still remains more robust than sGLP-IB.
Table 10: Results of our methods on a Raspberry Pi-4B device with four models on CIFAR-10. The values in parentheses indicate the percentage reduction in latency.
Latency(s) VGGNet-16 ResNet-56 ResNet-110 GoogleNet
Original 26.7 4.6 7.7 11.2
sGLP-IB 7.2(73%) 2.2(52%) 3.2(58%) 3.8(66%)
sTLP-IB 7.2(73%) 2.2(52%) 3.3(57%) 3.8(66%)
Implementation on Resource-constrained Devices
In addition to accuracy, memory size, and FLOPs, we also evaluate the latency we reduced using our method on resource-constrained devices. We use a Raspberry Pi 4B device. The results are presented in Table 10. We select the highest-accuracy models of our sGLP-IB and sTLP-IB methods for each of the four architectures listed in Table 1 of main paper. The results show that our method significantly reduces latency up to 73% while enabling the deployment of complex neural models on the resource-constrained devices.
Resource-constrained Devices
Our example experimental prototype Raspberry Pi 4B for resource-constrained devices is shown in Figure 3.